if a church steeple casts a shadow 112ft.long, and at the same time a 7.0ft. post cast a shadow 6.0 ft.long. How high is the church steeple
Answer by checkley77(12844)

7/6=x/112
6x=7*112
6x=784
x=784/6
x=130 2/3 ft. ans. for the height of the church steeple.